Ticket Number: HEA-NOC/971103-01 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <==> Ebone
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 971031 14:00 UTC
Ticket Closed: 19981028 12:29:33 UTC Problem End: 971103 15:30 UTC
All routes to Ebone were being filtered out because of a configuration
problem on Heuston.
Outbound traffic for European sites rerouted through Ten-34 or MCI;
incoming traffic was not affected.
Reconfigured Heuston and cleared the BGP session with Ebone.
Time to Fix:
3d 1h 30m approx.